# Can LLMs Handle WebShell Detection? Overcoming Detection Challenges with Behavioral Function-Aware Framework

### v2

@@ 4.2 The Effectiveness of Context-Aware Code Extraction (RQ2)
For the smaller model, Qwen-2.5-3B, critical regions significantly improve performance over the full source baseline. At $\tau=100$, the F1 score increases from 84.37% to 90.91% (+6.54%), with precision increasing from 78.03% to 86.71% (+8.68%) and recall increasing from 91.84% to 95.54% (+3.70%). This improvement validates the focus on behaviorally relevant code snippets, which reduces irrelevant context and sharpen the focus of the model. However, as the context length ($\tau$) increases, performance decreases - F1 drops to 88.17% at $\tau=300$, likely due to the model’s limited ability to handle extended context, which introduces noise that degrades accuracy.
@@ 4.3 The Effectiveness of WBFP for In-Context Learning (RQ3)
We evaluated the effectiveness of WBFP for ICL demonstration selection using Qwen-2.5-3B and GPT-4. This evaluation builds on the Context-Aware Code Extraction with the hybrid strategy. Four demonstration selection strategies were compared: Random Selection (Random), Source Code Semantic Similarity (SC-Sim), WBFP with Equal Weights (WBFP-Eq), and WBFP with Function-Level Weights (WBFP-Wt). Results are detailed in Tables 7 and 8 in Appendix D.
@@
Random selection makes ICL performance much worse by adding irrelevant examples, which makes both models much less effective. For Qwen-2.5-3B, the F1 score drops to 60.83% under the hybrid strategy with a context length of $\tau=100$, which is a 23.54% reduction from the no-ICL baseline of 84.37%. This is mainly because there is a large drop in precision from 78.03% to 46.33%, although recall remains strong at 88.53%. Similarly, for GPT-4 with $\tau=300$, the F1 score drops to 76.22%, a 16.24% decrease from the No-ICL baseline of 92.46%, driven by a drop in precision from 100.00% to 65.58%. These results indicate that Random Selection does not provide contextually relevant demonstrations, making it ineffective for WebShell detection.
The SC-Sim approach, which relies on semantic similarity calculated over entire source code samples, also doesn’t work well. For Qwen-2.5-3B, SC-Sim achieves an F1 score of 84.36%, which is almost the same as the No-ICL baseline (84.37%), with a precision of 78.57% and a recall of 91.08%. For GPT-4, it achieves an F1 score of 96.32%, with perfect precision, but a reduced recall of 92.90% compared to the No-ICL baseline’s 85.98%. This limited performance is likely due to the dominance of behaviorally irrelevant code segments in the similarity calculation, which dilutes the focus on critical behavioral patterns essential for accurate WebShell identification.
@@
For both models, WBFP-Wt consistently outperforms in accuracy, precision, recall, and F1 score, demonstrating its robustness and adaptability for improving ICL in WebShell detection tasks. Specifically, for Qwen-2.5-3B, WBFP-Wt achieves an F1 score of 93.69%, which is 9.33% higher than SC-Sim and 1.28% higher than WBFP-Eq. This is because WBFP-Wt achieves a precision of 88.64% (compared to 78.57% for SC-Sim and 86.39% for WBFP-Eq) and a near perfect recall of 99.36%. By focusing on the important parts of WebShell behavior, WBFP-Wt makes up for the fact that the smaller model doesn’t understand as much. This results in demonstrations that closely match the desired behavioral profiles, improving both precision and recall.
@@ 7 Conclusion
This paper presents a novel approach to WebShell detection using LLMs, addressing the unique challenges of applying these models to cybersecurity. We evaluated seven LLMs, including GPT-4 and LLaMA 3.1 70B, against traditional machine learning and deep learning methods using a dataset of 26.59K PHP scripts. Our analysis revealed that larger LLMs achieve near-perfect precision but lower recall, while smaller models exhibit high recall but poor precision, both of which underperform state-of-the-art methods such as GAT. To overcome these limitations, we proposed the BFAD framework, which integrates a critical function filter, context-aware code extraction, and WBFP. BFAD significantly improved LLM performance, with GPT-4 achieving an F1 of 99.35%, surpassing traditional benchmarks, and smaller models such as Qwen-2.5-0.5B improving by 51.23 percentage points to an F1 of 82.67%. As the first systematic exploration of LLMs for WebShell detection, this work not only demonstrates their potential, but also provides actionable solutions to their contextual and behavioral challenges, paving the way for future advances in code security.

### v3

@@ 4.2 The Effectiveness of Context-Aware Code Extraction (RQ2)
For the smaller model, Qwen-2.5-3B, critical regions significantly improve performance over the full source baseline. At $\tau=100$, the F1 score increases from 84.37% to 90.91% (+6.54%), with precision increasing from 78.03% to 86.71% (+8.68%) and recall increasing from 91.84% to 95.54% (+3.70%). This improvement validates the focus on behaviorally relevant code snippets, which reduces irrelevant context and sharpens the focus of the model. However, as the context length ($\tau$) increases, performance decreases - F1 drops to 88.17% at $\tau=300$, likely due to the model’s limited ability to handle extended context, which introduces noise that degrades accuracy.
@@ 4.3 The Effectiveness of WBFP for In-Context Learning (RQ3)
We evaluated the effectiveness of WBFP for ICL demonstration selection using Qwen-2.5-3B and GPT-4. This evaluation builds on the optimal hybrid strategy from RQ2. Four demonstration selection strategies were compared: Random Selection (Random), Source Code Semantic Similarity (SC-Sim), WBFP with Equal Weights (WBFP-Eq), and WBFP with Function-Level Weights (WBFP-Wt). Results are detailed in Tables 7 and 8 in Appendix D.
@@
Random selection makes ICL performance much worse by adding irrelevant examples, which makes both models much less effective. For Qwen-2.5-3B, the F1 score drops to 60.83% under the hybrid strategy with a context length of $\tau=100$, which is a 30.14% reduction from the no-ICL baseline of 90.97%. This is mainly because there is a large drop in precision from 89.02% to 46.33%, although recall remains strong at 88.53%. Similarly, for GPT-4 with $\tau=300$, the F1 score drops to 76.22%, a 22.16% decrease from the No-ICL baseline of 98.38%, driven by a drop in precision from 100.00% to 65.58%. These results indicate that Random Selection does not provide contextually relevant demonstrations, making it ineffective for WebShell detection.
The SC-Sim approach, which relies on semantic similarity calculated over entire source code samples, also doesn’t work well. For Qwen-2.5-3B, SC-Sim achieves an F1 score of 84.36%, a notable decrease from the No-ICL baseline of 90.97%, with a precision of 78.57% and a recall of 91.08%. For GPT-4, it achieves an F1 score of 96.32%, with perfect precision, but a reduced recall of 92.90% compared to the No-ICL baseline’s 96.82%. This limited performance is likely due to the dominance of behaviorally irrelevant code segments in the similarity calculation, which dilutes the focus on critical behavioral patterns essential for accurate WebShell identification.
@@
For both models, WBFP-Wt consistently outperforms in accuracy, precision, recall, and F1 score, demonstrating its robustness and adaptability for improving ICL in WebShell detection tasks. Specifically, for Qwen-2.5-3B, WBFP-Wt achieves an F1 score of 93.69%, which is 9.33% higher than SC-Sim and 1.28% higher than WBFP-Eq. This is because WBFP-Wt achieves a precision of 88.64% (compared to 78.57% for SC-Sim and 86.39% for WBFP-Eq) and a near-perfect recall of 99.36%. By focusing on the important parts of WebShell behavior, WBFP-Wt makes up for the fact that the smaller model doesn’t understand as much. This results in demonstrations that closely match the desired behavioral profiles, improving both precision and recall.
@@ 7 Conclusion
This paper presents a novel approach to WebShell detection using LLMs, addressing the unique challenges of applying these models to cybersecurity. We evaluated seven LLMs against traditional machine learning and deep learning methods using a dataset of 26.59K PHP scripts. Our analysis revealed that larger LLMs achieve near-perfect precision but lower recall, while smaller models exhibit high recall but poor precision, both of which underperform state-of-the-art methods such as GAT. To overcome these limitations, we proposed the BFAD framework, which integrates a critical function filter, context-aware code extraction, and WBFP. BFAD significantly improved LLM performance, with GPT-4 achieving an F1 of 99.35%, surpassing traditional benchmarks, and smaller models such as Qwen-2.5-0.5B improving by 51.23 percentage points to an F1 of 82.67%. As the first systematic exploration of LLMs for WebShell detection, this work provides actionable solutions to their contextual and behavioral challenges, paving the way for future advances in code security.